OpenStudy (anonymous):

What is the simplified form of (4x^2+20x+25)/(2x+5)

OpenStudy (anonymous):

try factorising 4x^2+20x+25 first

OpenStudy (anonymous):

how you cant the only thing their all divisable by is 1

OpenStudy (anonymous):

well you can 4x^2 +20x+25 = (2x+5)(2x+5) = (2x+5)^2

OpenStudy (anonymous):

oh foiling

OpenStudy (anonymous):

is the restriction of x positive or negitive?

OpenStudy (lgbasallote):

the restriction of x is just the value of x that will make 2x + 5 = 0

OpenStudy (anonymous):

so its negitive 5/2?

OpenStudy (lgbasallote):

yup that's the restricction of x

OpenStudy (anonymous):

check the attachement

OpenStudy (anonymous):

^ i don't understand what you did there @sami-21 , By factorising, we get [(2x+5)^2]/(2x+5) = 2x+5 which is the simplified form

OpenStudy (anonymous):

i just distributed the denominator you can get the same by taking LCM

OpenStudy (anonymous):

sorry i missed x with the 20 :(